HUBBLE SPACE TELESCOPE DAILY REPORT #2935 PERIOD COVERED: 0000Z (UTC) 08/13/01 - 0000Z (UTC) 08/14/01 Daily Status Report as of 226/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8907 (Spectroscopic Flats C10) The Space Telescope Imaging Spectrograph (CCD) was used to obtain CCD flats in the spectroscopic mode. There were no reported problems. 1.2 Completed Three Sets of WF/PC-2 8828 (Cycle 9 Supplemental Darks pt3/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed with no reported problems. 1.3 Completed Two Sets of STIS/CCD 9077 (Survey of the LMC Planetary Nebulae) The Space telescope Imaging Spectrograph (CCD) was used to perform a snapshot survey of all known LMC planetary nebulae {PNe} in order to study the co-evolution of the nebulae and their central stars, and to probe the chemical enrichment history of the LMC. There were no reported problems. 1.4 Completed Five Sets of WF/PC-2 9244 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The observations completed with no anomalous activity. 1.5 Completed Two Sets of STIS/CCD 8864 (CCD Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks for the CCD. The proposal completed nominally. 1.6 Completed Two Sets of WF/PC-2 8935 (Cycle 10 Standard Darks) The WF/PC-2 was used to obtain dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. The proposal completed with no reported problems. 1.7 Completed Seven Sets of STIS/CCD 8808 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted, parallel observations as part of a POMS test proposal. As described in 2.1 and HSTAR 8302, the acquisition for the final iteration of this proposal defaulted to fine lock backup on one FGS only, possibly affecting one observation. Otherwise, the observations completed with no further anomalous activity. 1.8 Completed Three Sets of WF/PC-2 9042 (An Archive To Detect The Progenitors Of Massive, Core-Collapse Supernovae) The WF/PC-s was used to search for supernovae which have massive star progenitors. The already extensive HST archive and high-resolution ground-based images of galaxies within ~20 Mpc enables us to resolve and quantify their individual bright stellar content. As massive, evolved stars are the most luminous single objects in a galaxy, the progenitors of core-collapse supernovae should be directly detectable on pre-explosion images. One Type II progenitor has been observed this year, and the investigators have proposed a short, companion WFPC2 proposal to confirm this candidate and identify a second. The observations completed nominally. 1.9 Completed STIS/CCD 8865 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.10 Completed STIS/CCD 8909 (CCD Dispersion Solutions) The Space Telescope Imaging Spectrograph (CCD) was used to perform wavelength dispersion solutions that are determined on a yearly basis as part of a long-term monitoring program and to obtain deep engineering wavecals for all CCD gratings at several wavelength centers. The proposal completed with no reported problems. 1.11 Completed Four Sets of WF/PC-2 8935 (Cycle 10 Standard Darks) The WF/PC-2 was used to obtain dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. As described in HSTAR 8301 and 2.1, the acquisition for the second iteration of this proposal defaulted to fine lock backup on one FGS only, possibly affecting one observation. Otherwise, the proposal completed with no other reported problems. 1.12 Completed STIS/CCD/MA1 8956 (Pre-SM Calibration) The Space Telescope Imaging Spectrograph (CCD and MA1) was used in an exercise to prepare for the SMOV-3B program. Prior to SMOV3B, one iteration of proposal 8960 {Jitter Test} for WD 2126+734 will be performed. Also one iteration of 8961 {End of BEA Test} for each new flux standard {WD 2126+734, WD 0320-539, WD 0710+741, and WD 0214+568} needed to completely cover all possible Bright Earth Avoidance {BEA} dates will be exercised. Finally, one extra standard is added {WD 1407-475} to fill the gap around June when WD 0214+568 is too close to the Sun. This iteration of the proposal completed with no reported problems. 1.13 Completed Three Sets of STIS/MA1/MA2 8843 (Cycle 9 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to perform the routine monitoring of the MAMA detector dark noise. The proposal completed nominally. 1.14 Completed WF/PC-2 8939 (Cycle 10 Internal Monitor) The WF/PC-2 was used to calibrate the internal monitor, to be run weekly to monitor the health of the cameras. No problems were encountered. 1.15 Completed Two Sets of WF/PC-2/STIS/MA1 8675 (The Massive Star Content of NGC 6822) The WF/PC-2 and the Space Telescope Imaging Spectrograph (MA1) were used to characterize the young, coeval stellar population discovered with WFPC2 multiband imaging in a NGC6822 star-forming region, with follow-up spectroscopy of the massive star candidates. The proposal completed with no reported problems. 1.16 Completed Three Sets of WF/PC-2 8941 (Cycle 10 UV Earthflats) The WF/PC-2 was used to monitor flat field stability by obtaining sequences of earth streak flats to improve the quality of pipeline flat fields for the WFPC2 UV filter set. The proposal had no problems. 1.17 Completed STIS/CCD/MA1 9051 (Identifying Damped Lyman-alpha Galaxies at z~1) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to look for samped Lyman-alpha absorption systems that contain the bulk of the neutral gas in the Universe in the redshift range z = 0.5 - 5 ,yet the nature of the galaxies responsible for the absorption is not well understood. Only recently have observers found more than a handful of damped absorbers at redshifts z < 1.5. Using the FIRST Bright Quasar Survey {FBQS}, with over a 1000 quasars, the proposers have undertaken a survey to build a complete picture of he nature of the galaxies responsible for damped Lyman-alpha absorption systems at z~1 and to double the sample size at this redshift. No problems were encountered. 1.18 Completed STIS/CCD 9165 (A Damped LyAlpha System in a Close Separation Lens) The Space Telescope Imaging Spectrograph (CCD) observed damped LyAlpha absorption systems that hold key information about galaxy formation and chemical evolution of the universe. It is proposed to conduct spatially resolved absorption line spectroscopy in the only presently known DLA system with two independent lines of sight. This object was discovered recently by us in the course of an ongoing HST snapshot survey as a double QSO at z=1.57 with an image separation of only 0.63 arcsec, almost certainly due to gravitational lensing. The ground-based composite optical spectrum shows an exceptionally strong Mg II+ Fe II absorption system at z=0.93, which clearly indicates the presence of a damped LyAlpha absorber {DLA} in the line of sight. The observations ended nominally. 1.19 Completed FGS/1 9169 (An Interferometric Harvest of Double Degenerates) Fine Guidance Sensor #1R was used to observe the white dwarf mass and age distributions that hold clues to the star formation history of our Galaxy and the age of the disk. No problems were reported. 1.20 Completed WF/PC-2 9057 (Host Galaxies of Obscured QSOs Identified by 2MASS) The WF/PC-2 was used to perform a snapshot survey of red QSOs discovered in The Two Micron All Sky Survey {2MASS} to investigate the detailed properties of their host galaxies. This large, possibly dominant, population of QSOs in the local universe has been previously overlooked because reddening by {intrinsic} obscuration along our line of sight causes their colors to be too red for identification by traditional "UV- excess" techniques. Their near-IR colors are similar to PG- type {UV-excess} QSOs, but it is far from certain whether they are indeed from the same parent population or represent a completely new class of QSO. As described in 2.1 and HSTAR 8302, the acquisition for this proposal defaulted to fine lock backup on one FGS only, possibly affecting two observations. Otherwise, there were no other reported problems. 1.21 Completed STIS/CCD 9285 (POMS Test Proposal: STIS Non-scripted Parallel Proposal Continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. As described in 2.1 and HSTAR 8302, the acquisition for this proposal defaulted to fine lock backup on one FGS only, possibly affecting one observation. Otherwise, the observations were completed as planned, and no further anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 12 Successful: 12 Per HSTAR 8301, the acquisition at 225/152216Z defaulted to fine lock backup on FGS-1 only. The proposal detailed in 1.11 may have been affected. The acquisition at 225/223920Z also defaulted to fine lock backup, this time on FGS-2 when the scan step limit was exceeded on FGS-1. The proposals described in 1.7, 1.20 and 1.21 may have been affected. Scheduled Re-acquisitions: 4 Successful: 4 2.2 FHST Updates: Scheduled: 33 Successful: 33 2.3 Operations Notes: Using ROP SR-1A, the SSR EDAC error counter was cleared three times. The ephemeris table was uplinked at 226/0203Z per ROP DF-07A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW